Two lanes of a motorway are being resurfaced.Each lane is 4 metres wide.The speed limit through the road works is 80 km/h.It takes 15 minutes to drive through the road works at this speed.Calculate the area being resurfaced
1
Expert's answer
2012-01-10T11:04:51-0500
As it takes 15 minutes ( = 0.25 hour ) to drive through the road with a speed of 80 km/h, then the length of the road works interval is L = 80*0.25 = 20 km. We know that the width of a lane is 4 metres ( = 0.04 km), so the area of a lane is A = 20*0.04 = 0.8 square kilometers. There are two resurfaced lanes so the result area is 0.8*2 = 1.6 km².